# Dynamic instabilities and turbulence of merged rotating Bose-Einstein condensates

### v1

@@ III Numerical Results
To study the merging process, we numerically solve the dynamical Gross-Pitaevskii (GP) equation (1) using the split-Step Crank-Nicolson method [46, 47, 48, 49]. Our simulations are performed on a spatial grid of size $512\times 512$ with space steps $dx=dy=0.05$ and time step $dt=10^{-4}$ to ensure convergence and accuracy. Similar to our previous work [50], the initial ground state consisting of two separate condensates is prepared using the imaginary time propagation scheme. Subsequently, we allow the merging of the condensates using the protocol discussed in the previous section, handled numerically using the real-time scheme for both rotating and non-rotating cases.
@@ III.1 Dynamical instabilities and turbulence of the merged condensate without rotation
Figure 1: Snapshots of the condensate density at (a) $t=0$, and at different instances for merging times $t_{m}=4$ [(b) - (e)] and $t_{m}=6$ [(f)-(j)]. (b) shows the formation of interference fringes alongside a central soliton, (c)-(d) shows the formation of vortex-antivortex pairs, and (e) represents the snapshot after the vortex pair has annihilated. (f) indicates the formation of interference fringes, (g) soliton with significant bending, (h)-(i) formation of vortex pairs, and (j) subsequent annihilation. The anti-vortices and vortices are marked with cross ($\times$) and circle ($\circ$) symbols, respectively.
@@
To begin with, we consider the merging dynamics of the condensate without any rotation. In this case, we prepare the initial state with a sufficient spatial separation between the two condensates to avoid any overlap between them. For a systematic investigation, we consider different merging times, viz., $t_{m}=4$, $6$, $10$, and $20$ between the condensates, and subsequently characterize their ensuing instabilities and turbulence states. Here, we have excluded the cases with lower merging times, such as $t_{m}<4$, because, for those events, the resulting condensate starts coming out from the single-well confinement due to the presence of uncontrolled large excitations induced in them. This situation prevents the condensate from exhibiting systematic energy spectra scalings, even over short intervals. We choose $t_{m}$ based on the rich dynamical instabilities inherited by the merged condensate, including the appearance of topological defects like vortices and their subsequent dynamics leading to the turbulent state.
@@
In Fig. 1, we show the evolution of the condensate density profiles at different instants of time for merging time $t_{m}=4$ and $t_{m}=6$ in the top [Figs. 1(a)-1(e)] and bottom panels [Figs. 1(f)-1(j)], respectively. For $t_{m}=4$, we notice the appearance of many intricate interference fringes, particularly a centralized dark soliton [panel(b)], which remains sustained in the newly formed single harmonic trap. However, as we follow up the dynamics for a longer duration, we find that the solitonic state does not remain stable, and they quickly get disintegrated into the combination of vortex-antivortex [panel (c)] pair as a result of the Snake instability [51, 52, 53]. Later, the vortex pair annihilates, generating compressible sound waves, as shown in panel (d). Additionally, compared to the density profiles observed at a slower merging time, we also observe the presence of interference fringes for the merging time $t_{m}=6$ [see Figs. 1(e)-(h)]. However, in this case, the soliton gets generated immediately after complete merging that exhibits significant bending [panel (f)], followed by a relatively quicker breakdown into vortex-antivortex pairs [panel (g)]. This particular feature triggers faster annihilation in the merged condensate compared to that of the $t_{m}=4$ case. For instance, for $t_{m}=6$ the condensates show a vortex-less state as a result of the vortex-antivortex annihilation even at $t\sim 400$, which was $t\sim 1800$ for $t_{m}=4$.
@@
Our analysis reveals that the merging time significantly affects the overall dynamics and structures of the merged condensate as the condensate transitions from the double well to the single well. To effectively capture this behaviour, Fig. 2 presents the temporal evolution of the condensate’s mean root-mean-squared radius ($\langle r\rangle_{rms}$) for various merging times.
@@
Next, we calculate the instantaneous forcing to understand the role of different competing factors involved in the instabilities of the merged condensate during various merging times. This forcing, being the time derivative of the space-averaged density spectra, reflects the appearance of sharp discontinuities in the condensate density at a specific time. Additionally, a finite jump in the force indicates the emergence of soliton formation [26]. Figure 3(a) illustrates the time evolution of the forcing for different merging times. It shows a decrease in forcing peak amplitude with longer merging times. This could potentially indicate phase discontinuities caused by the presence of solitons, which subsequently become less pronounced for larger merging times. The forcing peaks also decrease over time for a given merging time as the solitons eventually decay. Furthermore, the temporal evolution of the force exhibits an oscillation with a frequency of $2\omega$ which can be attributed to the presence of the harmonic trap with frequency $\omega$. These oscillations in the forcing value can be further related to the solitons bouncing off the trap boundary and propagating in the opposite $x$-direction.
@@
The phase profiles in Figs 3(b)-(e) highlight the variation in soliton profiles with respect to merging time. For faster mergers, we observe a larger number of solitons with sharper phase jumps [cf. Fig. 3(b)]. In contrast, slower mergers result in fewer solitons with smoother phase jumps [cf. Figs. 3(c)-(e)]. The soliton velocity depends on the phase difference across the soliton [33]. A larger phase jump indicates a narrow and deep soliton moving slowly, while a smoother phase gradient signifies a wider and shallower soliton moving closer to the speed of sound.
@@
To further analyze the mechanism responsible for the decay of solitons into vortex pairs, we investigate the temporal profile of the angular momentum and vortex number for $t_{m}=4$ and $6$, as depicted in Fig. 4. An early increase in angular momentum suggests the formation of point vortices due to the breaking of solitons in the non-rotating condensate. We then observe a subsequent decrease, indicating the annihilation of vortex-antivortex pairs. For longer merging times, the angular momentum peak appears earlier and decreases more rapidly, suggesting earlier soliton decay into vortex pairs and their subsequent annihilation.
One may witness strong evidence of soliton decay for different merging times in the vortex numbers illustrated in Fig. 4, where populations of vortices and anti-vortices appear to be equal for both merging times. The oscillations in the vortex number can be attributed to density oscillations, which are more pronounced for faster mergers ($t_{m}=4,6$).
Furthermore, evidence suggests the formation of dark neutral lumps (vortexoniums) from the coalescence of vortices and anti-vortices [51, 54, 55]. These vortex pairs undergo a cyclical process of forming lumps and returning to vortices until the vortexonium dissipates into compressible excitations. For longer merging times, the soliton becomes wide enough to reach the sound velocity before complete condensate merging. This results in dissipation without exhibiting snake instability, explaining the absence of vortex pairs in such cases.
@@
Next, we analyze the incompressible and compressible kinetic energies, which offer insights into the detailed nature of the fluctuations generated during the merging process, thereby quantifying the different characteristics of the merged condensate. The presence of point vortices contributes to the incompressible kinetic energy, whose evolution follows a similar trend as that of angular momentum and vortex numbers. The compressible kinetic energy also exhibits an early peak due to the trap merger generating sound waves, which eventually diminish and settle as the merged condensate evolves. Notably, the compressible kinetic energy is comparable to the incompressible kinetic energy due to the smaller number of vortices.
The Onsager energy, $E_{\mathrm{O}}$, i.e., the energy per vortex, is also computed and displayed in Fig. 5 for various merging times. The Onsager energy indicates that vortices formed through shorter mergers are more energetic and undergo oscillations consistent with a frequency $2.0\omega$.
Figures 6(a)-(d) illustrate the temporal evolution of the various energy components, such as the expectation value of the potential energy ($E_{pot}$), interaction energy ($E_{int}$), kinetic energy ($E_{kin}$), rotational energy ($E_{rot}$) for different merging time $t_{m}=4$, $6$, $10$, and $20$, respectively.
@@
We notice that the kinetic energy decays to zero for longer merging times ($t_{m}=10$, $20$), indicating that the merger is quasi-static for these $t_{m}$. Further, we find that the total energy gets significantly reduced as merging time increases, indicating a decrease in the excitation energy. This excitation energy depends on the depth of dark solitons, which are shallower for the slower merging cases, as shown in Fig. 6. A point to note is that all these time-varying quantities display oscillations with a frequency of $2.0\omega$ due to the nature of harmonic confinement. However, these oscillations get reduced when the merging reaches the quasi-static regime where the merger is slow enough not to instigate any dynamical process. These mergers appear to be the same as adiabatic mergers [26].
@@
Figure 7: Time-averaged incompressible kinetic energy spectra for the non-rotating case, for merging time $t_{m}=4$ in the time range (a) $t=4$ to $t=20$ (b) $t=20$ to $t=40$ and for merging time $t_{m}=6$ in the time range (c) $t=6$ to $t=20$ and (d) $t=20$ to $t=40$. Both merging cases initially show $k^{-5/3}$ scaling and later a $k^{-1}$ scaling.
Figure 8: Snapshots of the condensate density at (a) $t=0$ and at different instances during the condensate merger for merging times $t_{m}=6$ [(b) - (e)] and $t_{m}=10$ [(f)-(j)]. (b) Off-axis collisions forming weaker bent soliton structures, (c) soliton vortex interactions, (d)-(e) formation of single harmonically confined condensate with a vortex lattice, (f)-(g) shows off-axis collision with soliton-vortex interaction, and (h)-(j) appearance of the vortex lattice in the merged condensate.
@@
Next, to understand the detailed nature of the turbulence for different merging times, in Fig. 7, we present the incompressible kinetic energy spectra for various merging times. We observe that the merging condensate, except for the $t_{m}=4$, $6$ cases, does not exhibit any scaling behaviour in the inertial range. Additionally, the merged condensate immediately after merging ($t_{m}=4$, $6$) displays a $k^{-5/3}$ scaling for length scales smaller than the inter-vortex distance, $\ell_{0}$. This differs from the case where turbulence appears in a single condensate, as reported in Ref.  [50]. This behaviour can be attributed to the dominance of single vortex dynamics and the lack of sufficient vortices to display a clear Kolmogorov-like cascade of energy. However, at later times (significantly longer than the merging time), the condensate displays $k^{-1}$ scaling at regions smaller than the inter-vortex distance, indicating a Vinen-like cascade [56, 57, 50]. These scaling behaviours exhibited by the merged condensate are generally short-lived due to the annihilation of vortices in the absence of external mechanisms to generate new vortices, such as rotation. These features are also evident in the absence of a forcing peak at smaller $k$ values, indicating a lack of a proper mechanism to sustain the turbulent fluctuations in the merged condensate.
@@ III.2 Effect of rotation on the dynamics and turbulence state of the merged condensate
Having a comprehensive insight into the merger of two non-rotating condensates, we now turn our attention to investigating the effect of rotation on the merging dynamics of the condensate, starting with a low frequency of ($\Omega=0.45$). Similar to the non-rotating case for this case as well we achieve the merger by moving the traps and their rotating axes towards each other along the $x$-axis and simultaneously increasing their frequencies at the same rate, resulting in a final merged condensate rotating at $\Omega=0.9$.
In Figs. 8, we present the snapshots of the condensate at different instants of time for $t_{m}=6$ [8(a)-(e)] and $t_{m}=10$ [8(f)-(j)]. We notice that the condensates undergo an off-axis merger, deviating from the purely along-the-$x$-axis behaviour seen in the non-rotating case. This off-axis nature, arising from the movement of the rotation axis and trap towards the center of the simulation domain, becomes more pronounced for larger initial separations and shorter merging times ($t_{m}=6,10$). In such cases, the condensates can potentially spiral into each other, which strongly influences the nature of the soliton that appears as results of the instability, leading to distortions and orientation at an angle to the $x$-axis due to the presence of rotation. Soliton-vortex interactions also become more probable for slower mergers (see Fig. 8 bottom panel), but compared to the situation as discussed in [33], these interactions do not affect the vortex numbers or angular momenta in presence of the rotation.
@@
Next, we show the temporal evolution of the forcing of the merged condensate in Fig. 9(a) for different $t_{m}$, which follows a similar pattern to that of the non-rotating case but has a significantly smaller forcing amplitude due to the presence of shallower solitons that exhibit significant bending. Notably, the amplitude decays much slower when compared to the non-rotating case and persists for longer times due to the presence of rotational effects. The phase profiles in Figs 9(b)-(e) illustrate the nature of the soliton at the instant when merging takes place for $t_{m}=4$, $6$, $10$, and $20$ respectively for the rotating condensate. For small merging times ($t_{m}=4$, $6$), we notice the presence of a large number of solitons, which are quite evident from the discontinuous phase jumps [cf. Fig. 9(b,c)]. Conversely, for $t_{m}=10$, $20$, the phase profile is smoother, indicating the presence of fewer solitons and the emergence of vortices [cf. Figs. 9(d)-(e)]. Similar to the non-rotating case, we can relate the large velocity jump to the high velocity of the soliton, while the smooth phase corresponds to the soliton moving with the sound velocity [33].
@@
To characterize the effect of the rotation frequency on the force amplitude, which in turn affects the structure and dynamics of the soliton. In Fig. 10, we depict the variation of the force amplitude ($\tilde{f}_{amp}$) with the angular frequency ($\Omega$) of the rotation for different $t_{m}$. We find that for small merging times ($t_{m}=4$, $6$), the force amplitude decreases with an increase in $\Omega$However, for longer merging times ($t_{m}=10$, $20$), the force amplitude is close to zero and independent of $\Omega$.
Figure 11: Evolution of vortex ($N_{\mathrm{v}}$) and anti-vortex ($N_{\mathrm{av}}$) numbers within the Thomas-Fermi radius of the condensate ($R_{\mathrm{TF}}$) for (a) $t_{m}=4$ (b) $t_{m}=6$ (c) $t_{m}=10$ (d) $t_{m}=20$ for rotating case with $\Omega=0.45$.
The angular momentum shows a significantly higher vortex count at shorter merging times. This phenomenon arises from the formation and subsequent decay of solitons into vortex-antivortex pairs, contributing to the angular momentum in cases where $t_{m}=4,6$. For longer merging times, the likelihood of soliton formation diminishes, resulting in a decrease in the contribution of solitons to both angular momentum and vortex numbers. Additionally, the rotational effect causes the vortex population to surpass the anti-vortex population. Figure 11 illustrates the evolution of vortex numbers over time. Notably, the presence of rotation leads to significantly more vortices than anti-vortices, and the existence of vortex pairs confirms soliton decay. For faster mergers in Figure 11(a), the vortex number oscillates at $2\omega$ due to density oscillations. However, we witness a significant reduction of this behaviour for slower mergers in Figs. 11(c)-(d), following a trend similar to the previous case.
@@
The incompressible kinetic energy evolution exhibits a decreasing trend concerning merging time. In contrast to the non-rotating scenario, the incompressible energy does not quickly settle down. Instead, it persists alongside the angular momentum profiles, indicating a continuous production of vortices. Examining the compressible kinetic energy reveals a trend similar to that of the non-rotating case. However, for longer merging times ($t_{m}=20$), there is a persistent and constant compressible energy, while in the non-rotating case, the compressible energy approaches zero at these times. This discrepancy is primarily due to the presence of rotational forcing, which not only generates more vortices but also enhances the production of compressible sound waves. The Onsager energy profile for different merging times in the rotating case is quite different, as shown in Fig. 12. In the rotating case, vortices are not only formed via decaying solitons but also due to the presence of rotation. From Fig.12 one can see that vortices formed via rotation have more Onsager energy when compared to the ones formed via solitons. The lowest Onsager energy for $t_{m}=4$ (black line) can be attributed to the significant increase in the number of vortices formed due to snake instability, that lowers the overall Onsager energy in the condensate despite the presence of vortices formed due to rotation.
@@
In Fig. 13(a)-(d), we show the various energy components, $E_{pot}$, $E_{int}$, $E_{kin}$, $E_{rot}$ for different merging times, i.e., $t_{m}=4$, $6$, $10$, and $20$, respectively, keeping the rotation frequency fixed at $\Omega=0.45$.

### v3

@@ III Numerical Results
To study the merging process, we numerically solve the dynamical Gross-Pitaevskii (GP) equation (1) using the split-Step Crank-Nicolson methodMuruganandam and Adhikari (2009); Vudragović et al. (2012); Young-S et al. (2017); Kumar et al. (2019). Our simulations are performed on a spatial grid of size $512\times 512$ with space steps $dx=dy=0.05$ and time step $dt=10^{-4}$ to ensure convergence and accuracy. Similar to our previous workSivakumar et al. (2024), the initial ground state consisting of two separate condensates is prepared using the imaginary time propagation scheme. Subsequently, we allow the merging of the condensates using the protocol discussed in the previous section, handled numerically using the real-time scheme for both rotating and non-rotating cases.
@@ III.1 Dynamical instabilities and turbulence of the merged condensate without rotation
To begin with, we consider the merging dynamics of the condensate without any rotation. For this case, we prepare the initial state of two identical condensates with a random phase which helped us to take account of the effect of quantum noise present in actual experimental setups. For a systematic investigation, we consider different merging times, viz. $t_{m}=4$, $6$, $10$, and $20$ between the condensates and subsequently characterize their ensuing instabilities and turbulence states. To avoid the escape of resulting condensate quickly from the single-well confinement due to uncontrolled large excitations we excluded such cases of short merging times $(t_{m}<4)$ This prevents the condensate from exhibiting consistent energy spectra scaling, even for short durations. We selected $t_{m}$ based on the complex dynamical instabilities observed in the merged condensate, including the formation of topological defects like vortices and their subsequent chaotic dynamics leading to a turbulent state.
Figure 3: Three-dimensional visualizations of the condensate density surface at (a) $t=0$ [cf. Fig. 2(a)], (b) $t=2$, and various time instances, (c) $t=4$, (d) $t=12$, (e) $t=400$ and (f) $t=1800$, corresponding to the merging time $t_{m}=4$ [cf. Fig. 2(b) - 2(e)].
@@
Figure 2 illustrates the evolution of condensate density profiles for merging times $t_{m}=4$ and $t_{m}=6$. The top panels (a)-(e) depict the case for $t_{m}=4$, while the bottom panels (f)-(j) show the case for $t_{m}=6$. For a more comprehensive visualization of the merging dynamics, in Fig. 3, we display the surface plots of condensate densities for $t_{m}=4$ corresponding to the time points depicted in Figs. 2(a)-2(e), including an intermediate time $t=2$ in Fig. 3(b). For $t_{m}=4$, we notice the appearance of many intricate interference fringes, particularly a centralized dark soliton [panel(b)], which remains sustained in the newly formed single harmonic trap. However, as we follow up the dynamics for a longer duration, we find that the solitonic state does not remain stable, and quickly disintegrates into the combination of the vortex-antivortex [panel (c)] pair as a result of the Snake instabilityVerma, Rapol, and Nath (2017); Theocharis et al. (2003); Mamaev, Saffman, and Zozulya (1996). Later, the vortex pair annihilates, generating compressible sound waves, as shown in panel (d). Additionally, compared to the density profiles observed at a slower merging time, we also observe the presence of interference fringes for the merging time $t_{m}=6$ [see Figs. 2(e)-(h)]. However, in this case, the soliton that gets generated immediately after complete merging exhibits significant bending [panel (f)], followed by a relatively quicker breakdown into vortex-antivortex pairs [panel (g)]. This particular feature triggers faster annihilation in the merged condensate compared to that of the $t_{m}=4$ case. For example, for $t_{m}=6$, the condensates show a vortexless state as a result of the vortex-antivortex annihilation even at $t\sim 400$, which was $t\sim 1800$ for $t_{m}=4$. The dark-soliton fringes initially formed symmetrically during the merger. However, as the merger progresses, the central soliton bends to the right and oscillates by reflecting off the trap boundaries, consistent with the equations of motion for dark solitons, as shown in Ref.Frantzeskakis, 2010.
@@
Our analysis reveals that the merging time significantly affects the overall dynamics and structures of the merged condensate as the condensate transitions from the double well to the single well. To effectively capture this behaviour, Fig. 4 presents the temporal evolution of the mean root mean squared radius of the condensate ($\langle r\rangle_{rms}$) for various merging times.
@@
Next, we calculate the instantaneous forcing to understand the role of different competing factors involved in the instabilities of the merged condensate during various merging times. This forcing, which is the time derivative of the space-averaged density spectra, reflects the appearance of sharp discontinuities in the condensate density at a specific time. Additionally, a finite jump in force indicates the emergence of soliton formationBuchmann, Nikolopoulos, and Lambropoulos (2009). In this case, the forcing term lies within the interval $\sim[-10^{-4},10^{-4}]$ with the standard deviation of approximately $10^{-7}$ for $t_{m}=4$ and $t_{m}=6$, and that of the order of $10^{-8}$ for $t_{m}=10$ and $t_{m}=20$. Figure 5(a) illustrates the time evolution of the forcing for different merging times. It shows a decrease in the forcing peak amplitude with longer merging times. This could potentially indicate phase discontinuities caused by the presence of solitons, which subsequently become less pronounced for larger merging times. The forcing peaks also decrease over time for a given merging time as the solitons eventually decay. Furthermore, the temporal evolution of the force exhibits an oscillation with a frequency of $2\omega$ which can be attributed to the presence of the harmonic trap with frequency $\omega$. These oscillations in the forcing value can be further related to the solitons that bounce off the trap boundary and propagate in the opposite $x$ direction. The phase profiles in Figs 5(b)-(e) highlight the variation in soliton profiles with respect to merging time. For faster mergers, we observe a larger number of solitons with sharper phase jumps [see Fig. 5(b)]. In contrast, slower mergers result in fewer solitons with smoother phase jumps [see Figs. 5(c)-(e)]. The soliton velocity depends on the phase difference across the solitonKanai, Guo, and Tsubota (2019). A larger phase jump indicates a narrow and deep soliton moving slowly, while a smoother phase gradient signifies a wider and shallower soliton moving closer to the speed of sound.
@@
To further analyze the mechanism responsible for the decay of solitons into vortex pairs, we investigate the temporal profile of the vortex number for $t_{m}=4$ and $6$, as depicted in Fig. 6. We witness strong evidence of soliton decay for different merging times in the vortex numbers illustrated in Fig. 6, where populations of vortices and anti-vortices appear to be equal for both merging times. The oscillations in the vortex number can be attributed to density oscillations, which are more pronounced for faster mergers ($t_{m}=4$, $6$). Furthermore, evidence suggests the formation of dark neutral lumps (vortexoniums) from the coalescence of vortices and anti-vorticesVerma, Rapol, and Nath (2017); Smirnov and Smirnov (2015); Groszek et al. (2016). These vortex pairs undergo a cyclical process of forming lumps and returning to vortices until the vortexonium dissipates into compressible excitations. For longer merging times, the soliton becomes wide enough to reach the sound velocity before complete condensate merging. This results in dissipation without exhibiting snake instability, which explains the absence of vortex pairs in such cases.
@@
Next, we analyze the spatial distribution of the incompressible (not shown) and compressible kinetic energies [see Fig. A.1 in Appendix A] for $t_{m}=4$, which offer insights into the detailed nature of the fluctuations generated during the merging process and subsequent instabilities leading to the turbulent state, thereby quantifying the different characteristics of the merged condensate. The presence of point vortices contributes to the incompressible kinetic energy, whose evolution follows a similar trend as that of vortex numbers. The compressible kinetic energy also exhibits an early peak [see Fig. A.1(a)-(c)] due to the trap merger generating sound waves, which eventually diminish and settle as the merged condensate evolves [see Fig. A.1(d)]. Notably, the compressible kinetic energy is comparable to the incompressible kinetic energy because of the smaller number of vortices.
The Onsager energy, $E_{\mathrm{O}}$, that is, the energy per vortex, is also calculated and displayed in Fig. 7 for various merging times. The Onsager energy indicates that vortices formed through shorter mergers are more energetic and undergo oscillations consistent with a frequency $2.0\omega$.
Figures 8(a)-(d) illustrate the temporal evolution of the various energy components, such as the expectation value of the potential energy ($E_{pot}$), interaction energy ($E_{int}$), kinetic energy ($E_{kin}$), rotational energy ($E_{rot}$) for different merging times $t_{m}=4$, $6$, $10$, and $20$, respectively.
@@
Figure 9: Time-averaged incompressible kinetic energy spectra for the non-rotating case, for merging time $t_{m}=4$ in the time range (a) $t=4$ to $t=14$ (b) $t=14$ to $t=40$ and for merging time $t_{m}=6$ in the time range (c) $t=6$ to $t=14$ and (d) $t=14$ to $t=40$. Both merging cases initially show $k^{-5/3}$ scaling and later a $k^{-1}$ scaling
We notice that the kinetic energy decays to zero for longer merging times ($t_{m}=10$, $20$), indicating that the merger is quasi-static for these $t_{m}$. Further, we find that the total energy gets significantly reduced as merging time increases, indicating a decrease in the excitation energy. This excitation energy depends on the depth of dark solitons, which are shallower for the slower merging cases, as shown in Fig. 8. A point to note is that all these time-varying quantities display oscillations with a frequency of $2.0\omega$ due to the nature of harmonic confinement. However, these oscillations get reduced when the merging reaches the quasi-static regime where the merger is slow enough not to instigate any dynamical process. These mergers appear to be the same as adiabatic mergersBuchmann, Nikolopoulos, and Lambropoulos (2009). Interestingly, the oscillations between potential and kinetic energy suggest that our merger process resembles the turbulence generation mechanism described in the recent work of Middleton-Spencer et al.Middleton-Spencer et al. (2023). In their study, an oscillating external potential nucleates vortex and anti-vortex pairs through solitons, which ultimately annihilate to produce density waves. Our system exhibits analogous instabilities following the merging, ultimately culminating in a turbulent state.
@@
Next, to understand the detailed nature of the turbulence for different merging times, in Fig. 9, we present the incompressible kinetic energy spectra for various merging times. We observe that the merging condensate, except for the $t_{m}=4$, $6$ cases, does not exhibit any scaling behaviour in the inertial range. Additionally, the merged condensate immediately after merging ($t_{m}=4$, $6$) displays a $k^{-5/3}$ scaling for length scales smaller than the inter-vortex distance, $\ell_{0}$. This differs from the case where turbulence appears in a single condensate, as reported in Ref. Sivakumar et al., 2024. This behaviour can be attributed to the dominance of single vortex dynamics and the lack of sufficient vortices to display a clear Kolmogorov-like cascade of energy. However, at later times (significantly longer than the merging time), the condensate displays $k^{-1}$ scaling at regions smaller than the inter-vortex distance, indicating a Vinen-like cascadeMarino et al. (2021); Cidrim et al. (2017); Sivakumar et al. (2024). These scaling behaviours exhibited by the merged condensate are generally short-lived due to the annihilation of vortices in the absence of external mechanisms to generate new vortices, such as rotation. These features are also evident in the absence of a forcing peak at smaller $k$ values, indicating a lack of a proper mechanism to sustain turbulent fluctuations in the merged condensate. To confirm the nature of the turbulence, we performed the merging simulation for a larger grid size. In this case, the incompressible spectra show similar scaling behaviour where both mergers initially display $k^{-5/3}$ and later $k^{-1}$ scalings (see Fig. 9). Even at larger grid sizes, the presence of $k^{-5/3}$ scaling is not in the range $k<2\pi/\ell_{0}$, indicating that a Kolmogorov cascade is not present.
@@ III.2 Effect of rotation on the dynamics and turbulence state of the merged condensate
Having a comprehensive insight into the merger of two non-rotating condensates, we now turn our attention to investigating the effect of rotation on the merging dynamics of the condensate, starting with a frequency of ($\Omega_{0}=0.45$).
Figure 10: Snapshots of the condensate density for the condensates rotating initially at $\Omega_{0}=0.45$ at different instants of time: (a) $t=0$ and at different instances during the condensate merger for merging times $t_{m}=6$ [(b)-(e)] and $t_{m}=10$ [(f)-(j)]. (b) Off-axis collisions forming weaker bent soliton structures, (c) soliton vortex interactions, (d)-(e) formation of single harmonically confined condensate with a vortex lattice, (f)-(g) show off-axis collision with soliton-vortex interaction, and (h)-(j) appearance of the vortex lattice in the merged condensate.
In the rotating merger, we move the traps and their rotating axes towards each other along the $x$-axis while increasing their rotation frequencies at the same rate, resulting in a final merged condensate rotating at $\Omega_{f}=2\Omega_{0}=0.9$.
In Figs. 10, we present the snapshots of the condensate at different instants of time for $t_{m}=6$ [Figs. 10(a)-(e)] and $t_{m}=10$ [Figs. 10(f)-(j)]. We notice that the condensates undergo an off-axis merger, deviating from the purely along-the-$x$-axis behaviour seen in the non-rotating case. This off-axis nature, arising from the movement of the rotation axis and trap towards the center of the simulation domain, becomes more pronounced for larger initial separations and shorter merging times ($t_{m}=6,10$). In such cases, the condensates can potentially spiral into each other, which strongly influences the nature of the soliton that appears as results of the instability, leading to distortions and orientation at an angle to the $x$-axis due to the presence of rotation. Soliton-vortex interactions also become more probable for slower mergers (see Fig. 10 bottom panel), but compared to the situation as discussed in Ref. Kanai, Guo, and Tsubota, 2019, these interactions do not affect the vortex numbers or angular momenta in presence of the rotation.
Figure 11: Three-dimensional visualizations of the condensate density surface at (a) $t=0$ [cf. Fig. 10(a)], (b) $t=2$, and various time instances, (c) $t=4$, (d) $t=6$, (e) $t=20$ and (f) $t=1000$, corresponding to the merging time $t_{m}=6$ [cf. Fig. 10(b) - 10(e)].
Figure 11 provides a comprehensive visualization of the surface plots of condensate densities for rotational merger with $t_{m}=6$ corresponding to the time points depicted in Figs. 10(a)-10(e), including an intermediate time $t=2$ in Fig. 11(b).
@@
Next, we show the temporal evolution of the forcing of the merged condensate in Fig. 12(a) for different $t_{m}$, which follows a similar pattern to that of the non-rotating case but has a significantly smaller forcing amplitude due to the presence of shallower solitons that exhibit significant bending. In this case, the forcing term lies within the interval $\sim[-10^{-5},10^{-5}]$ with the standard deviation of approximately $10^{-8}$ for $t_{m}=4$ and $t_{m}=6$, and that of the order of $10^{-9}$ for $t_{m}=10$ and $t_{m}=20$. Notably, the amplitude decays much slower when compared to the non-rotating case and persists for longer times due to the presence of rotational effects. The long correlation for the forcing can be further confirmed through the slow decay of the compressible energy peak as illustrated in Fig. A.2 (see Appendix A). The phase profiles in Figs. 12(b)-(e) illustrate the nature of the soliton at the instant when merging takes place for $t_{m}=4$, $6$, $10$, and $20$ respectively for the rotating condensate. For small merging times ($t_{m}=4$, $6$), we notice the presence of a large number of solitons, which are quite evident from the discontinuous phase jumps [see Fig. 12(b),(c)]. Conversely, for $t_{m}=10$, $20$, the phase profile is smoother, indicating the presence of fewer solitons and the emergence of vortices [see Figs. 12(d)-(e)]. Similar to the non-rotating case, we can relate the large velocity jump to the high velocity of the soliton, while the smooth phase corresponds to the soliton moving with the sound velocityKanai, Guo, and Tsubota (2019).
@@
To characterize the effect of the rotation frequency on the force amplitude, which in turn affects the structure and dynamics of the soliton. In Fig. 13, we depict the variation of the force amplitude ($\tilde{f}_{amp}$) with the angular frequency ($\Omega_{0}$) of the rotation for different $t_{m}$. We find that for small merging times ($t_{m}=4$, $6$), the force amplitude decreases with an increase in $\Omega_{0}$However, for longer merging times ($t_{m}=10$, $20$), the force amplitude is close to zero and independent of $\Omega_{0}$.
Figure 14: Evolution of vortex ($N_{\mathrm{v}}$) and anti-vortex ($N_{\mathrm{av}}$) numbers within the Thomas-Fermi radius of the condensate ($R_{\mathrm{TF}}$) for (a) $t_{m}=4$ (b) $t_{m}=6$ (c) $t_{m}=10$ (d) $t_{m}=20$ for rotating case with $\Omega_{0}=0.45$.
There is a significantly higher vortex count at shorter merging times. This phenomenon arises from the formation and subsequent decay of solitons into vortex-antivortex pairs, contributing to the vortex-population in cases where $t_{m}=4,6$. For longer merging times, the likelihood of soliton formation diminishes, resulting in a decrease in the contribution of soliton decay to the vortex population. Additionally, we find that the rotational effect causes the vortex population to surpass the anti-vortex population which is quite evident from the evolution of vortex numbers as shown in Figure 14. For faster mergers in Fig. 14(a), the vortex number oscillates at $2\omega$ due to density oscillations. However, we witness a significant reduction of this behaviour for slower mergers in Figs. 14(c)-(d), following a trend similar to the previous case.
@@
The incompressible kinetic energy evolution exhibits a decreasing trend concerning merging time. In contrast to the non-rotating scenario, the incompressible energy does not quickly settle down. Instead, it persists indicating a continuous production of vortices. Examining the compressible kinetic energy reveals a trend similar to that of the non-rotating case. However, for longer merging times ($t_{m}=20$), there is a persistent and constant compressible energy, while in the non-rotating case, the compressible energy approaches zero at these times. This discrepancy is primarily due to the presence of rotational forcing, which not only generates more vortices but also enhances the production of compressible sound waves. The Onsager energy profile for different merging times in the rotating case is quite different, as shown in Fig. 15. In the rotating case, vortices are not only formed via decaying solitons but also due to the presence of rotation. From Fig. 15 one can see that vortices formed via rotation have more Onsager energy when compared to the ones formed via solitons. The lowest Onsager energy for $t_{m}=4$ (black line) can be attributed to the significant increase in the number of vortices formed due to snake instability, which lowers the overall Onsager energy in the condensate despite the presence of vortices formed due to rotation.
@@
In Fig. 16(a)-(d), we show the various energy components, $E_{pot}$, $E_{int}$, $E_{kin}$, $E_{rot}$ for different merging times, i.e., $t_{m}=4$, $6$, $10$, and $20$, respectively, keeping the rotation frequency fixed at $\Omega_{0}=0.45$.